Point M(12, 18) is given in a rectangular plane. Which is the distance of this point from the pole?

A. 12.63
B. 18.36
C. 20.46
D. 21.63
1
Expert's answer
2013-09-04T08:53:32-0400
Since the pole in the polar coordinate system is the origin (the point (0,0)) in Cartesian coordinate system), the distance from the point M to the pole is the distance from the point M to the origin. To calculate the distance d from (0,0) to (12,18), use the formula d = sqrt (12 * 12 +18 * 18) = sqrt (468) = 21.63.
Answer: D. 21.63